Circular No. 7041

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998em IN IC 4837A J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova (B ∼ 16.5) by Roberto Antezana on a T-Max 400 film taken by L. E. González with the Maksutov telescope at Cerro El Roble on Oct. 26.96 UT. SN 1998em is located at α = 19^{h}15^{m}12^{s}.69, δ = −54^{o}06′49′′.7 (equinox 2000.0), which is 26′′ west and 67′′ north of the spiral galaxy IC 4837A. The new object is not present on a film taken on Oct. 13.96 (B > 19). SN 1998em was confirmed on direct CCD images taken by I. Perez on Oct. 27.95 using the 1.0-m Swope telescope at Las Campanas Observatory. A spectrum was obtained by W. Kunkel, Carnegie Institution of Washington (CIW), using the du Pont telescope (+ Modular Spectrograph) at Las Campanas. M. Phillips, CIW, reports: “The spectrum of the supernova exhibits strong, broad Si II 635.5-nm absorption consistent with a type-Ia event at maximum, or perhaps a few days before. Narrow Galactic interstellar Na I D lines with a combined equivalent width of 0.054 nm are observed near zero velocity. A second set of Na I D absorption lines with an equivalent width of 0.11 nm and presumably due to gas in the host galaxy IC 4837A is present at a redshift of 0.0091.”

1ES 2005–489 R. Remillard reports for the ASM instrument teams at the Massachu-

setts Insitute of Technology and Goddard Space Flight Center: “The RXTE All Sky Monitor has detected increased x-ray flux from the BL Lac object 1ES 2005–489 = PKS 2005–489 (α = 20^{h}09^{m}25^{s}.39, δ = −48^{o}49′53′′.8, equinox 2000.0). During the interval Oct. 11–28, the flux has been in the range of 4 to 10 mCrab at 2–12 keV, which is a factor of 2–3 above its typical intensity during 1996–1998. Observations at other wavelengths are strongly encouraged.”

NOVA SCORPII 1998 W. Liller, Viña del Mar, Chile, reports the following magnitudes for N

Sco 1998 as measured with a broadband-V system (CCD + minus-infrared filter). Note that the first magnitude has been revised slightly from what appeared on IAUC 7037: Oct. 23.0029 UT, 7.71; 23.9983, 8.08; 25.0006, 8.34; 25.9993, 8.45.

Additional visual magnitude estimates by W. G. Dillon, Missouri City, TX: Oct. 25.05 UT, 9.2; 26.03, 9.4; 27.03, 9.6.

1998 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green